OpenStudy (anonymous):

in store A the price of mayonnaise went down from $9.66 to $6.25. in store B, it went from $6.41to $1.45. in store C it went from $4.29 to $3.66 . what was the average decrease?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

What was this price decrease for each store?

OpenStudy (anonymous):

idk it goes by A,B,C

OpenStudy (uri):

Start with A What is 9.66-6.25?

OpenStudy (anonymous):

ommmg 0.41

OpenStudy (sleepyjess):

Store A - price decrease = 9.66-6.25 Store B - price decrease = 6.41-1.45 Store C - price decrease = 4.29-3.66

OpenStudy (uri):

...Use a calculator..

OpenStudy (anonymous):

i am getting a headahche of not getting it, it takes to long to keep restartinf

OpenStudy (anonymous):

an i did

OpenStudy (anonymous):

i need something to eat brb

OpenStudy (uri):

@mcnabkt Well you should rest...because I think your headache won't let you learn,yeah do that.

OpenStudy (anonymous):

i just need to finish these questions

OpenStudy (sleepyjess):

So, what is the price decrease for store A?

OpenStudy (anonymous):

3.41

OpenStudy (sleepyjess):

Good, store B?

OpenStudy (anonymous):

B.4.96 C . 0.63

OpenStudy (sleepyjess):

Good! Now add all 3 of those price decreases.

OpenStudy (anonymous):

9?

OpenStudy (sleepyjess):

Yep, now divide 9/3

OpenStudy (anonymous):

3 lol

OpenStudy (sleepyjess):

Yep!

OpenStudy (anonymous):

so the answer is 3?

OpenStudy (sleepyjess):

Yes

OpenStudy (uri):

Nice gurl ^